A MAN IS FIVE YEARS OLDER THAN HIS WIFE. THIRTEEN YEARS AGO HE WAS TWICE HER AGE. HOW OLD IS THE MAN?
PLEASE HELP WITH THIS I AM HAVING A HARD TIME SETTING THIS ONE UP.
THANKS IN ADVANCE

Answer by stanbon(48546)

A MAN IS FIVE YEARS OLDER THAN HIS WIFE. THIRTEEN YEARS AGO HE WAS TWICE HER AGE. HOW OLD IS THE MAN?
--------------------------
Now DATA:
wife's age = x yrs.
man's age = x+5 yrs.
------------------------
13 yrs. ago DATA:
wife was x-13 yrs
man was x-8 yrs
------------------------
EQUATION:
x-8 = 2(x-13)
x-8 = 2x-26
x = 18 (wife's age now)
x+5 = 23 (man's age now)
=============================
Cheers,
Stan H.